OpenStudy (anonymous):

Simplify 3(x - 2) + 5x 1. 8x - 2 2. 8x - 6 3. 6x My bet's on 1, but...here you go @Dwade03

OpenStudy (anonymous):

nope 2

OpenStudy (anonymous):

3(x-2) = 3x-6 +5x= 8x-6

OpenStudy (anonymous):

8x-6

OpenStudy (anonymous):

Ah, thanks. Might as well post the next to here. Simplify 6 - 3(4 - 3x) 1. -3x - 6 2. 9x - 6 3. 9x + 18

OpenStudy (anonymous):

2. 9x-6

OpenStudy (anonymous):

And finally... Simplify 2(x + 4) + 3(x - 4) 1. 5x - 4 2. 5x 3. 5x + 4 I'm guessing that it's 2. Not because the correct answers are coincidentally the second one, but because +4 and -4 cancel each other out, and it's down to adding. I could be wrong though.

OpenStudy (anonymous):

its 1.

OpenStudy (anonymous):

Yeah, @Sephichief, you need you distribute the parenthesis and then add like-terms.

OpenStudy (anonymous):

2(x+4)=2x+8 then add 3(x-4) =3x -12 so add both together 2x+8+3x-12= 5x-4

OpenStudy (anonymous):

@friedch1cken So you mean my first assumption was correct, or...?

OpenStudy (anonymous):

No, I meant @Dwade03 was correct when he said it's 1 :D

OpenStudy (anonymous):

you have to distribute it first and then add

OpenStudy (anonymous):

You two just saved me from having to do all of my previous assignments over again.

OpenStudy (anonymous):

A shame I can't give best response to both of you.

OpenStudy (anonymous):

what you mean

OpenStudy (anonymous):

Don't worry about it. Thanks!

OpenStudy (anonymous):

lol k :)